# Improved Streaming Algorithm for Fair $k$-Center Clustering

### v1

@@ 4 Streaming Semi-structured Data Sets
In this section, we consider scenarios in which the data points are streamed following the demographic group order, i.e. all points belong to $S_{l}$ arrive before $S_{l+1}$, $\forall l\in[m-1]$. We provide an improved algorithm achieving a ratio 3 for $m=2$ that conforms to the state-of-art ratio for offline setting. Moreover, for general $m$, we show that the algorithm can be extended to achieve a ratio $4$ (as shown in Appendix).
Our key idea of the improved algorithm for $m=2$ is to obtain an extra point set $\Gamma_{sub}$ in the streaming stage, such that points of $\Gamma_{sub}$ can be used to replace the points of $\Gamma_{1}$ for the fairness constraint. In addition, to better suit $\Gamma_{sub}$, we compute a slightly improved independent center set $\Gamma_{l}^{\prime}$ instead of $\Gamma_{l}$ except for $\Gamma_{1}$. We first construct an independent set $\Gamma_{1}^{\prime}=\Gamma_{1}$ with $\lambda=2r^{*}$ upon the data stream of $S_{1}$. Then, upon the stream of $S_{2}$, we select points using different designate approaches depending on whether $|\Gamma_{1}^{\prime}|\leq k_{1}$ holds. The detailed algorithm is as in Alg. 2.
Then we prove that with only the representative points of $\Gamma_{1}^{\prime}\cup\Gamma_{2}^{\prime}$ and $\Gamma_{sub}$, our post-streaming algorithm can compute an approximation solution of ratio 3.
@@
Output: The sets of representative points.
2
3Set $\Gamma_{sub}=\emptyset$ and $\Gamma_{l}^{\prime}=\emptyset$ for $l=1,2$;
@@
7
8
9upon each arriving point $i\in S_{2}$ do
10
@@ 21         Set $\Gamma_{sub}\leftarrow\Gamma_{sub}\cup\left\{i\right\}$;
22
23
24
25
26Return $\Gamma_{1}^{\prime}$, $\Gamma_{2}^{\prime}$, and $\Gamma_{sub}$.
Algorithm 2 Improved streaming stage against semi-structural streams
@@
For $\Gamma_{1}^{\prime}$ and $\Gamma_{2}^{\prime}$ produced by the above algorithm, we have: (1) $|\Gamma_{1}^{\prime}|+|\Gamma_{2}^{\prime}|\leq k$; (2) $\Gamma_{2}^{\prime}\leq k_{2}$ always holds; (3) There exists a subset $\Gamma_{1}^{\prime\prime}\subseteq\Gamma_{1}^{\prime}$ and accordingly a subset $\Gamma^{\prime}_{sub}=\{\sigma(c)|c\in\Gamma_{1}^{\prime\prime}\}\subseteq\Gamma_{sub}$, such that $|\Gamma_{1}^{\prime}\setminus\Gamma_{1}^{\prime\prime}\cup\Gamma^{\prime}_{sub}\cap S_{1}|\leq k_{1}$ holds, where $\sigma(c)$ is a possible replacement of $c$ in $\Gamma_{sub}$, i.e. $d(c,\sigma(c))\leq r^{*}$.
@@
For Cond. (1), according to the aglorithm, any pair of two points $p,q\in\bigcup\Gamma_{l}^{\prime},l=1,2$, we have $d(p,q)>2r^{*}$. So every pair of points must appears in two different clusters in the optimal solution. Thus, $|\bigcup_{l=1}^{2}\Gamma_{l}^{\prime}|\leq k$ holds.
For Cond. (2), we have $|\Gamma_{2}^{\prime}|\leq k-|\Gamma_{1}^{\prime}|<k-k_{1}=k_{2}$ if $|\Gamma_{1}^{\prime}|>k_{1}$. Otherwise, as $\Gamma_{1}^{\prime}$ covers all points of the clusters centered at points of $S_{1}$ in the optimal solution within a radius $3r^{*}$. Then, the remaining points to be covered in the algorithm appears in only the optimal clusters centered at points of $S_{2}$, which are at most $k_{2}$ clusters. On the other hand, every pair of points in $\Gamma_{2}^{\prime}$ must appear in different optimal clusters. So $|\Gamma_{2}^{\prime}|\leq k_{2}$.
For Cond. (3), if $|\Gamma_{1}^{\prime}|\leq k_{1}$ holds after streaming then the proof is done. So we need only to prove the case for $|\Gamma_{1}^{\prime}|>k_{1}$. Clearly, there exists a subset of $\Gamma_{1}^{\prime}$ contains at least $|\Gamma_{1}^{\prime}|-k_{1}$ points, say $\Gamma_{1}^{\prime\prime}$, appearing in optimal clusters centered at points of $S_{2}$. According to Alg. 2, at least one point $j\in S_{2}$ for each point $i\in\Gamma_{1}^{\prime\prime}$ exists with $d(i,j)\leq r^{*}$, and is added as $\sigma(i)$ to $\Gamma_{sub}$ as a replacement point of $i$, collectively composing the set $\Gamma^{\prime}_{sub}$. So $|\Gamma_{1}^{\prime}\setminus\Gamma_{1}^{\prime\prime}\cup\Gamma^{\prime}_{sub}\cap S_{1}|\leq k_{1}$ holds. ∎
Following the above lemma, we can immediately obtain a 3-approximation solution by simply computing $C=\Gamma_{2}\cup\Gamma_{1}^{\prime}\setminus\Gamma_{1}^{\prime\prime}\cup\Gamma^{\prime}_{sub}$ as the desired center set.
@@ 5 Experimental Results
In this section, we conduct an empirical evaluation of our algorithms utilizing both simulated and real-world datasets, compared with three previous approximation algorithms that serve as baselines. All experiments are averaged over at least 20 iterations and conducted on a Linux machine equipped with 12th Gen Intel(R) Core(TM) i9-12900K CPU and 32 GB RAM by Python 3.8. A detailed description of the experimental setup is introduced below.
@@ 5.1 Experimental setting
Real-world datasets: We apply our algorithms to seven real-world datasets: Wholesale, Student, Bank, CreditCard, Adult, SushiA and CelebA, following the most recent related works (Jones, Nguyen, and Nguyen 2020; Chen et al. 2019). Consistent with previous studies, we utilize meaningful numerical features for clustering and incorporate selected binary categorical attributes to construct datasets with fair constraints across all datasets.
Simulated datasets: we provide two datasets that serve as complementary benchmarks: the simulated dataset provides an exact optimal solution, whereas the large-scale dataset includes more time for file I/O operations, making it ideal for measuring computational efficiency. First, the dataset (called SimuA) is used to assess the empirical approximation ratio of our algorithms and the baseline methods. Inspired by previous research (Kleindessner, Awasthi, and Morgenstern 2019), we used their method to construct a simulated dataset with a known optimal solution for the fair $k$-center problem. Secondly, we leverage the implementation from Chiplunkar, Kale, and Ramamoorthy (2020) to generate a 100G dataset (called SimuB), which allows us to evaluate the runtime performance of different algorithms.
Constraints Settings
For the simulated dataset, we configured the parameters to compare both the average and maximum values of the approximation ratios for these algorithms and to verify the approximation ratio of our algorithms. To ensure fair center selection, we aligned the number of required centers from each group with the proportional size of that group. Following the fairness principle of disparate impact as outlined by Feldman et al. (2015), we restricted the selection to $k_{l}$ data points from the $l$th group to serve as centers. We then evaluated the clustering quality by varying the number of $k$ across these datasets.
@@
Algorithms
Compared with existing streaming fair $k$-center clustering algorithms, this experiment includes an extended algorithm from the offline 3-approximation algorithm by Jones, Nguyen, and Nguyen (2020) (denoted as ExJones), a two-pass streaming 3-approximation algorithm by Chiplunkar, Kale, and Ramamoorthy (2020) (denoted as two-pass $3$-Approx), and a one-pass streaming 7-approximation algorithm (Lin, Guo, and Jia 2024) (denoted as $7$-Approx). Moreover, we propose two algorithms in this paper: 1) a one-pass 5-approximation algorithm (denoted as $5$-Approx) on the general metric; 2) a one-pass 3-approximation algorithm with $m=2$ (denoted as $3$-Approx) for semi-structured data streaming in our experiments.
Metrics
We use the cost metric, as defined in Sec. 2, to compare the quality of clustering across these datasets based on their average values. In addition, we adopted the second to measure runtime.
Dataset #Record #Dim.
@@ 5.2 Experimental analysis
In Fig. 1, we first report the empirical approximation ratios of the algorithms on a simulated dataset with known optimal solutions. This allows us to compare the experimental objective values of our algorithms and the baselines against the optimal objective. We then evaluate the clustering cost of these algorithms on real-world datasets. Finally, we report the runtime(s) performance on large-scale simulated datasets, focusing on file I/O operations in our algorithms compared to other baselines.
Approximation Factor
We compare algorithm performance by evaluating the relative solution ratio with the provided optimal solution on the simulated dataset. The ratio of the evaluation result can be called the empirical approximation ratio, and the maximum value represents the worst-case cost in all the experimental results. We run the code of constructing the simulated dataset shared by Kleindessner, Awasthi, and Morgenstern (2019) for their algorithm setting $m=2$, $k=100$ and varying the size of the simulated datasets $|S|$ from $2,000$ to $10,000$. For each size of the dataset, we perform $10$ runs on $10$ kinds of fairness with $10$ different constraint ratios.
We observe that all algorithms align well with their theoretical bounds presented in this paper and the baselines. Notably, the $3$-Approx algorithm demonstrates the best performance, even outperforming the two-pass 3-Approx algorithm (Chiplunkar, Kale, and Ramamoorthy 2020). This aligns with our theoretical results, as the algorithm provides a strong approximation guarantee. When the dataset size $|S|=2,000$, the maximal empirical approximation ratio of the 7-Approx algorithm achieves an approximation ratio greater than 5, validating the suitability of the dataset for testing these algorithms. Compared with the Jones’ method, our algorithm has a lower empirical approximation ratio, which demonstrates the advantage of our algorithm’s result. Consequently, these algorithms exhibit superior performance on ideal datasets, aligning well with their theoretical guarantees.
Dataset ExJones 77-Approx 55-Approx
33-Approx
33-Approx
(Two pass)
@@
Table 2: Cost comparison on the real-world datasets. (The underline highlights the best results of the general one-pass algorithms. In addition, the bold indicates the best result when semi-structured data streaming is included alongside the one-pass streaming algorithms.)
Comparison with Baselines
@@
In Tab. 2, we observe that $3$-Approx (Semi) outperforms most of the other one-pass algorithms, while the two-pass 3-approximation algorithm keeps the better performace than the one-pass streaming algorithms on the seven datasets. We attribute this to the fact that our algorithm is a modification of the original center selection method based on streaming. By combining streaming techniques with the structure of independent center set, it effectively identifies more meaningful center points. As a result, it achieves lower clustering costs than the other baselines under the one-pass streaming setting (shown as the underline value). We also observe that the results reported by $3$-Aprrox (two-pass) consistently outperform all the one-pass streaming algorithms. This can be explained that the two-pass 3-approximation algorithm has the advantage of refining the center set in the second pass, enabling it to identify more accurate center points.
@@
Streaming Runtime
Using the method from Chiplunkar, Kale, and Ramamoorthy (2020), we generate an artificial 100 GB dataset containing 4,000,000 items and 1,000 features. In Fig. 2, we evaluate the performance at run-time of our algorithms and baseline methods in this dataset. Our results show that $5$-Approx algorithm runs faster than Chiplunkar, Kale, and Ramamoorthy (2020) and Jones, Nguyen, and Nguyen (2020), which aligns with the theoretical expectations, as all of our algorithms are one-pass streaming methods. Among these methods, 7-Approx (Lin, Guo, and Jia 2024) achieves the fastest runtime, because it selects fewer points during the streaming process and omits post-processing steps, thereby reducing computational overhead.
@@ 6 Conclusion
In this paper, we first devise a one-pass streaming algorithm with an approximation ratio of $5$ and a memory complexity of $O(k\log n)$. We demonstrate that such approximation ratio and memory usage are optimal for the metric space. Observing the broad applications of semi-structured data streams, we present a 3-approximation for $m=2$ and a 4-approximation for general $m$. Lastly, extensive experiments were conducted to demonstrate the performance gain of our algorithms compared to baselines including the state-of-the-art method.

### v2

@@ 4 Streaming Semi-structured Datasets
In this section, we consider scenarios in which the data points are streamed following the demographic group order, i.e. all points belonging to $S_{l}$ arrive before $S_{l+1}$, $\forall l\in[m-1]$. We devise an algorithm for $m=2$, achieving a ratio of $3+\epsilon$ that almost matches the state-of-the-art ratio of $3$ for the offline setting. The algorithm can be extended to achieve a ratio $4$ for general $m$ as shown in the full version.
Our key idea of the improved algorithm for $m=2$ is to obtain an extra point set $\Gamma_{sub}$ in the streaming stage and use some of its points as centers for points in $\Gamma_{1}$. More precisely, we first construct an independent set $\Gamma_{1}^{\prime}=\Gamma_{1}$ with $\lambda=2r^{*}$ upon the data stream of $S_{1}$. Then, upon the stream of $S_{2}$, we select points using different designated approaches depending on whether $|\Gamma_{1}^{\prime}|\leq k_{1}$ holds. As the constructed independent center sets are slightly different, we adopt notation $\Gamma_{l}^{\prime}$ instead of $\Gamma_{l}$. The detailed algorithm is as in Alg. 2.
Then we prove that with only the representative points of $\Gamma_{1}^{\prime}\cup\Gamma_{2}^{\prime}$ and $\Gamma_{sub}$, our post-streaming algorithm can compute an approximation solution with ratio $3$.
@@
2Set $\Gamma_{sub}=\emptyset$ and $\Gamma_{l}^{\prime}=\emptyset$ for $l=1,2$;
@@
5
6
7upon each arriving point $i\in S_{2}$ do
@@ 18         Set $\Gamma_{sub}\leftarrow\Gamma_{sub}\cup\left\{i\right\}$;
19
20
21
Return $\Gamma_{1}^{\prime}$, $\Gamma_{2}^{\prime}$, and $\Gamma_{sub}$.
Algorithm 2 Streaming semi-structured data
@@
For $\Gamma_{1}^{\prime}$ and $\Gamma_{2}^{\prime}$ produced by Alg. 2, we have: (1) $|\Gamma_{1}^{\prime}|+|\Gamma_{2}^{\prime}|\leq k$; (2) $|\Gamma_{2}^{\prime}|\leq k_{2}$; (3) There exists a subset $\Gamma_{1}^{\prime\prime}\subseteq\Gamma_{1}^{\prime}$ and accordingly a subset $\Gamma^{\prime}_{sub}=\{\sigma(c)|c\in\Gamma_{1}^{\prime\prime}\}\subseteq\Gamma_{sub}$, such that $|\Gamma_{1}^{\prime}\setminus\Gamma_{1}^{\prime\prime}\cup\Gamma^{\prime}_{sub}\cap S_{1}|\leq k_{1}$ holds, where $\sigma(c)$ is the stored replacement of $c$ in $\Gamma_{sub}$ with $d(c,\sigma(c))\leq r^{*}$.
@@
For Cond. (1), according to the algorithm, we have $d(p,q)>2r^{*}$ for any pair of two points $p,q\in\bigcup\Gamma_{l}^{\prime},l=1,2$. So every pair of points of $\bigcup\Gamma_{l}^{\prime}$ must appear in two different clusters in the optimal solution similar to the proof of Lem 2. Thus, $|\bigcup_{l=1}^{2}\Gamma_{l}^{\prime}|\leq k$ holds.
For Cond. (2), if $|\Gamma_{1}^{\prime}|>k_{1}$, we have $|\Gamma_{2}^{\prime}|\leq k-|\Gamma_{1}^{\prime}|<k-k_{1}=k_{2}$ . Otherwise (i.e., $|\Gamma_{1}^{\prime}|\leq k_{1}$), as $\Gamma_{1}^{\prime}$ covers all points of the clusters centered at points of $S_{1}$ in the optimal solution within a radius $3r^{*}$. Then, the remaining points to be covered in the algorithm appear in only the optimal clusters centered at points of $S_{2}$, which are at most $k_{2}$ clusters. On the other hand, every pair of points in $\Gamma_{2}^{\prime}$ must appear in different optimal clusters. So $|\Gamma_{2}^{\prime}|\leq k_{2}$.
For Cond. (3), if $|\Gamma_{1}^{\prime}|\leq k_{1}$ holds after streaming then the proof is done. So we need only to prove the case for $|\Gamma_{1}^{\prime}|>k_{1}$. Clearly, there exists a subset of $\Gamma_{1}^{\prime}$ that contains at least $|\Gamma_{1}^{\prime}|-k_{1}$ points, say $\Gamma_{1}^{\prime\prime}$, appearing in optimal clusters centered at points of $S_{2}$. Clearly, there exists at least one point $j\in S_{2}$ for each point $i\in\Gamma_{1}^{\prime\prime}$ with $d(i,j)\leq r^{*}$. Moreover, according to Alg. 2, we add such point $\sigma(i)=j$ to $\Gamma_{sub}$ as a replacement point of $i$, collectively composing the set $\Gamma^{\prime}_{sub}$. That is, $\Gamma_{1}^{\prime}\setminus\Gamma_{1}^{\prime\prime}\cup\Gamma^{\prime}_{sub}$ feasibly covers all points of $\Gamma^{\prime}_{1}$ with radius $r^{*}$, and contains at most $k_{1}$ points of $\Gamma_{1}^{\prime}$. ∎
Following the above lemma, we can immediately obtain a 3-approximation solution by simply using $C=\Gamma^{\prime}_{2}\cup(\Gamma_{1}^{\prime}\setminus\Gamma_{1}^{\prime\prime}\cup\Gamma^{\prime}_{sub})$ as the desired center set.
@@ 5 Experimental Results
In this section, we conduct an empirical evaluation of our algorithms utilizing both simulated and real-world datasets, compared with three previous approximation algorithms that serve as baselines. All experiments are averaged over at least 20 iterations and are conducted on a Linux machine equipped with a 12th Gen Intel(R) Core(TM) i9-12900K CPU and 32 GB of RAM using Python 3.8. A detailed description of the experimental setup is provided in follows.
@@ 5.1 Experimental Setting
Real-world datasets. We apply our algorithms to seven real-world datasets: Wholesale (Cardoso 2013), Student (Cortez 2008), Bank (Moro et al. 2014), CreditCard (Yeh 2009), Adult (Becker and Kohavi 1996), SushiA and CelebA (Liu et al. 2015), following the most recent related works (Jones et al. 2020; Chen et al. 2019). Consistent with previous studies, we utilize meaningful numerical features for clustering and incorporate selected binary categorical attributes to construct datasets with fair constraints across all datasets.
Simulated datasets. We provide two datasets that serve as benchmarks: the simulated dataset provides an exact optimal solution, while the large-scale dataset incurs more file I/O overhead, making it ideal for measuring computational efficiency. First, the dataset (called SimuA) is used to assess the empirical approximation ratio of our algorithms and the baseline methods. Inspired by previous research (Kleindessner et al. 2019), we use their method to construct a simulated dataset with a known optimal solution for the fair $k$-center problem. Secondly, we leverage the implementation from Chiplunkar et al. (2020) to generate a 100 GB dataset (called SimuB), which allows us to evaluate the runtime performance of different algorithms.
Constraint Settings.
For the simulated dataset, we configured the parameters to compare both the average and maximum approximation ratios for these algorithms and to verify the approximation ratio of our algorithms. To ensure fair center selection, we aligned the number of required centers from each group to be proportional to the size of that group. Following the fairness principle of disparate impact as outlined by Feldman et al. (2015), we restricted the selection to $k_{l}$ data points from the $l$th group to serve as centers. We then evaluated the clustering quality by varying the number of $k$ clusters across these datasets.
@@
Algorithms.
In comparison with existing streaming fair $k$-center clustering algorithms, this experiment includes an extended algorithm from the offline 3-approximation algorithm by Jones et al. (2020) (denoted as ExJones), a two-pass streaming 3-approximation algorithm by Chiplunkar et al. (2020) (denoted as two-pass $3$-Approx), and a one-pass streaming 7-approximation algorithm (Lin et al. 2024) (denoted as $7$-Approx). Moreover, we propose two algorithms in this paper: 1) a one-pass 5-approximation algorithm (denoted as $5$-Approx); 2) a one-pass 3-approximation algorithm with $m=2$ (denoted as $3$-Approx) for semi-structured data streams in our experiments.
Metrics. To evaluate clustering quality, we use the cost metric defined in Sec. 2 and compare its average value across datasets. We also report runtime, measured in seconds.
Dataset #Records #Dim.
@@ 5.2 Experimental Analysis
In Fig. 1, we first report the empirical approximation ratios of the algorithms on a simulated dataset with known optimal solutions. This allows us to compare the experimental objective values of our algorithms and the baselines against the optimal objective. We then evaluate the clustering costs of these algorithms on real-world datasets. Finally, we report the runtime(s) performance on large-scale simulated datasets, focusing on file I/O operations in our algorithms compared to other baselines.
Approximation Factor.
We compare algorithm performance by evaluating the relative solution ratio against the optimal solution provided for the simulated dataset. The ratio of the evaluation result can be called the empirical approximation ratio, and the maximum value represents the worst-case cost in all the experimental results. We run the code of constructing the simulated dataset shared by Kleindessner et al. (2019) for their algorithm setting $m=2$, $k=100$ and varying the size of the simulated datasets $|S|$ from $2,000$ to $10,000$. For each size of the dataset, we perform $10$ runs on $10$ kinds of fairness with $10$ different constraint ratios.
Dataset ExJones 77-Approx 55-Approx 33-Approx (Semi) 33-Approx (Two pass)
@@
Table 2: Cost comparison on the real-world datasets. (The underline values highlight the best results of the general one-pass algorithms. In addition, the bold indicates the best result when semi-structured data streaming is included alongside the one-pass streaming algorithms.)
We observe that the practical performance of all algorithms aligns well with their theoretical bounds. Notably, the $3$-Approx (semi) demonstrates the best performance, even outperforming the two-pass 3-Approx algorithm (Chiplunkar et al. 2020). When the dataset size $|S|=2,000$, the maximal empirical approximation ratio of 7-Approx is greater than 5, validating the suitability of the dataset for testing these algorithms. Compared with the method of Jones et al., our algorithm achieves a lower empirical approximation ratio, demonstrating the advantage of our algorithm. Consequently, our algorithms exhibit superior performance on the datasets, conforming to the theoretical guarantees.
Comparison with Baselines.
@@
In Tab. 2, we observe that $3$-Approx (Semi) outperforms most of the other one-pass algorithms, while the two-pass $3$-approximation algorithm maintains better performance than the one-pass streaming algorithms on the seven datasets. We attribute this to the fact that our algorithm is a modification of the original center selection method based on streaming. By combining streaming techniques with the structure of an independent center set, it effectively identifies more meaningful center points. As a result, it achieves lower clustering costs than the other baselines under the one-pass streaming setting (shown as the underline values). We also observe that the results reported by (two-pass) $3$-Approx consistently outperform all the one-pass streaming algorithms. This can be explained by the fact that the two-pass 3-approximation algorithm has the advantage of refining the center set in the second pass, enabling it to identify more accurate center points.
@@
Streaming Runtime.
Using the method from Chiplunkar et al. (2020), we generate a 100 GB dataset containing 4,000,000 items and 1,000 features. In Fig. 2, we evaluate the runtime performance of our algorithms and baseline methods on this dataset. Our results show that the $5$-Approx algorithm runs faster than the algorithms of Chiplunkar et al. and Jones et al. (2020), which aligns with theoretical expectations, as all of our algorithms are one-pass streaming methods. Among these methods, 7-Approx (Lin et al. 2024) runs the fastest, because it selects fewer points during the streaming process and omits post-processing steps, thereby reducing computational overhead.
@@ 6 Conclusion
In this paper, we first devise a one-pass streaming algorithm with an approximation ratio of $5+\epsilon$ and a memory complexity of $O(k\log\alpha)$, where $\alpha$ is the aspect ratio. Motivated by the broad applications of semi-structured data streams, we then present an algorithm with a ratio of $(3+\epsilon)$ and the same memory complexity for the problem on semi-structured streams with $m=2$. This $(3+\epsilon)$-approximation can be directly used to solve the offline problem and achieve a slightly improved ratio of 3, matching the state-of-the-art ratio for offline fair $k$-center. Lastly, we conduct extensive experiments to demonstrate the performance gains of our algorithms compared to baselines, including state-of-the-art methods. In the future, we will strive to design both offline and streaming algorithms with better approximation ratios for fair $k$-center, $k$-median, and $k$-means in doubling metrics and Euclidean spaces.
Acknowledgments
This work is supported by National Natural Science Foundation of China (No. 12271098) and Key Project of the Natural Science Foundation of Fujian Province (No. 2025J02011).